Cindy is 5 years older than Tom. In 6 years, the sum of their ages will be 35 years. How old is each person? How old is each person now? Thanks for your help.
Answer by stanbon(75887)

Cindy is 5 years older than Tom. In 6 years, the sum of their ages will be 35 years. How old is each person? How old is each person now? Thanks for your help. ------
Equations:
t-c = 5
t+c = 35
----------
2t = 40
Tom = 20 years
-----
Cindy = 35-5 = 15 years
---------
Cheers,
Stan H.
----------